# Model X-ray : Detect Backdoored Models via Decision Boundary

### v1

@@
Abstract
Deep neural networks (DNNs) have revolutionized various industries, leading to the rise of Machine Learning as a Service (MLaaS). In this paradigm, well-trained models are typically deployed through APIs. However, DNNs are susceptible to backdoor attacks, which pose significant risks to their applications. This vulnerability necessitates a method for users to ascertain whether an API is compromised before usage. Although many backdoor detection methods have been developed, they often operate under the assumption that the defender has access to specific information such as details of the attack, soft predictions from the model API, and even the knowledge of the model parameters, limiting their practicality in MLaaS scenarios.
@@ Abstract
To address it, in this paper, we begin by presenting an intriguing observation: the decision boundary of the backdoored model exhibits a greater degree of closeness than that of the clean model. Simultaneously, if only one single label is infected, a larger portion of the regions will be dominated by the attacked label. Building upon this observation, we propose Model X-ray , a novel backdoor detection approach for MLaaS through the analysis of decision boundaries. Model X-ray can not only identify whether the target API is infected by backdoor attacks but also determine the target attacked label under the all-to-one attack strategy. Importantly, it accomplishes this solely by the hard prediction of clean inputs, regardless of any assumptions about attacks and prior knowledge of the training details of the model. Extensive experiments demonstrated that Model X-ray can be effective for MLaaS across diverse backdoor attacks, datasets, and architectures.
Machine Learning, ICML
@@
Figure 1: Comparison of the decision boundaries between the clean model and the backdoored model (taking BadNets (Gu et al., 2017) as an example, and the target label is “airplane”) on the CIFAR-10 dataset.
1 Introduction
Deep Neural Networks (DNNs) have exhibited remarkable success in various domains, such as computer vision (Krizhevsky et al., 2012) and speech recognition (Graves et al., 2013). Capitalizing on their exceptional performance, Machine Learning as a Service (MLaaS) has emerged as a significant sector within the AI industry. The robust growth of the MLaaS industry can also be attributed to major public cloud providers like Google, AWS, and Microsoft, who have played a crucial role in facilitating the deployment of MLaaS by offering computational power and scalability. To safeguard the intellectual property (IP) of these vendors, many services, such as ChatGPT, are made available through Application Programming Interfaces (APIs), which are often treated as black-box systems.
@@ 1 Introduction
Despite the remarkable success of DNNs, recent studies (Gu et al., 2017; Li et al., 2021a; Zeng et al., 2021; Wang et al., 2022; Liu et al., 2018; Doan et al., 2021; Bagdasaryan & Shmatikov, 2021) have unveiled a significant security vulnerability in DNNs. The vulnerability arises from backdoor attacks, which can contaminate DNNs, enabling them to operate normally on clean inputs but manipulate predictions whenever a specific pattern (i.e., “trigger”) is detected. Backdoor attacks primarily fall into two categories: data-poisoning attacks (such as BadNets (Gu et al., 2017), SSBA (Li et al., 2021a), Low Frequency (Zeng et al., 2021), and BPP (Wang et al., 2022)) and model-modification attacks (such as TrojanNN (Liu et al., 2018), LIRA (Doan et al., 2021), and Blind (Bagdasaryan & Shmatikov, 2021)). These attacks pose a substantial threat to MLaaS in safety-critical and security-sensitive applications, including but not limited to face recognition (Parkhi et al., 2015), biomedical diagnosis (Esteva et al., 2017), and autonomous driving (Redmon et al., 2016). Therefore, it is crucial for users to determine whether the API is compromised before utilizing it.
@@
To mitigate the threat of backdoor attacks, numerous detection methods are emerging to establish a comprehensive pipeline for backdoor defense. This pipeline can be applied at various stages, including the training, post-training, and deployment phases (refer to Fig. 2). As depicted in Table 1, backdoor detection (Chen et al., 2018; Tran et al., 2018; Gao et al., 2019; Zeng et al., 2021; Liu et al., 2023) during both the training and deployment phases typically necessitates access to training data or inference data, a requirement often impractical in MLaaS scenarios. Comparably, post-training backdoor detection is designed to evaluate whether a trained model has been compromised by malicious functions. However, current methods hold too strong assumptions that the defender has access to attack information, the soft prediction from the model API (Chen et al., 2019; Xu et al., 2021), and even require access to the knowledge of the model parameters (Wang et al., 2019; Liu et al., 2019; Fu et al., 2023), limiting their practicality in MLaaS scenarios we focus on.
@@
In this paper, our goal is to develop a backdoor detection approach for MLaaS, namely, empowering users to assess the security of a model API before usage. Specifically, the user has no access to the training dataset or any backdoored samples. Additionally, we are focusing on an exceedingly challenging scenario where users can only obtain predicted hard labels for clean inputs. Fortunately, recent work by (Somepalli et al., 2022) has demonstrated that we can visualize the model’s decision boundary solely using hard predictions. Leveraging this technique, we have identified a discernible distinction between the decision boundaries of the clean model and the backdoored model. As illustrated in Fig. 1, we use BadNets (Gu et al., 2017) as an example of backdoor attacks. We observe that the decision boundary of the backdoored model is more gathered compared to that of the clean model. Additionally, a significant portion of the decision boundary is controlled by the target label, encircling the clean region. Importantly, this phenomenon is applicable across various backdoor attacks on different datasets (see Fig. 4). As claimed in previous work (Wang et al., 2019), backdoor attacks build a shortcut between the trigger patterns and the target label, which we explain cause the above encircling phonemona. Besdies, trigger samples are more robust against distortions (Rajabi et al., 2023), which cause the large regions. In a nutshell, the visualized 2D decision boundary can be served as an illustration for these conjectures.
@@
Based on the intriguing phenomenon, we propose Model X-ray , a novel backdoor detection approach for MLaaS via the visualized decision boundary. Specifically, we designate two metrics to evaluate the degree of the closeness of the decision boundary: 1) Rényi Entropy (RE) (ren, ) calculated on the probability distribution of each prediction area and 2) Areas Dominated by triple samples (ATS), e.g., the total areas of “frog”, “ship”, and “dog” in the Fig. 1. Furthermore, if only one label is infected, we can determine the target label by the prediction of the largest area of the decision boundary, e.g., the target label is “airplane” in the right of Fig. 1. In a nutshell, Model X-ray can not only identify backdoored models but also determine the target attacked label under all-to-one attacks. Importantly, Model X-ray accomplishes this only by the predicted hard label of benign inputs, regardless of any assumptions about attacks such as the trigger patterns and training details. Extensive experiments demonstrate that Model X-ray performs better than current methods across various backdoor attacks, datasets, and model architectures. In addition, some ablation studies and discussions are also provided.
Figure 2: The pipeline of the backdoor defense.
@@
  • We summarize existing backdoor detection methods, categorizing them based on detection stage, defender capabilities, and detection capabilities (refer to Table 1). Additionally, we emphasize the limitations of these methods, particularly their impracticality in Machine Learning as a Service (MLaaS) scenarios.
@@
  • We present a noteworthy observation: there exists a distinction between clean models and backdoored models by the visualization of the decision boundary (Somepalli et al., 2022).
@@
  • We propose Model X-ray , which detects the backdoored model solely by predicted hard labels of clean inputs, regardless of any assumptions about backdoor attacks. Besides, Model X-ray can determine the target attacked label if the attack is all-to-one attack.
  • Extensive experiments demonstrate the effectiveness of Model X-ray across different backdoor attacks, datasets, and model architectures.
@@
[…]
Spectral Training
STRIP Deployment
FreqDetector Deployment
TeCo Deployment
Neural Cleanse Post-training
MNTD Post-training
MM-BD Post-training
[…]
Table 1: Qualitative comparison with current backdoor detection methods according to detection stage, defender capabilities, and detection capabilities. “”, “”, and “” represent unsatisfied, partially satisfied, and fully satisfied, respectively.

### v2

# Model X-ray : Detecting Backdoored Models via Decision Boundary

@@
Abstract.
Backdoor attacks pose a significant security vulnerability for deep neural networks (DNNs), enabling them to operate normally on clean inputs but manipulate predictions when specific trigger patterns occur. In this paper, we consider a practical post-training scenario backdoor defense, where the defender aims to evaluate whether a trained model has been compromised by backdoor attacks. Currently, post-training backdoor detection approaches often operate under the assumption that the defender has knowledge of the attack information, logit output from the model, and knowledge of the model parameters, limiting their implementation in practical scenarios.
@@ Abstract.
In contrast, our approach functions as a lightweight diagnostic scanning tool offering interpretability and visualization. By accessing the model to obtain hard labels, we construct decision boundaries within the convex combination of three samples. We present an intriguing observation of two phenomena in backdoored models: a noticeable shrinking of areas dominated by clean samples and a significant increase in the surrounding areas dominated by target labels. Leveraging this observation, we propose Model X-ray, a novel backdoor detection approach based on the analysis of illustrated two-dimensional (2D) decision boundaries. Our approach includes two strategies focused on the decision areas dominated by clean samples and the concentration of label distribution, and it can not only identify whether the target model is infected but also determine the target attacked label under the all-to-one attack strategy. Importantly, it accomplishes this solely by the predicted hard labels of clean inputs, regardless of any assumptions about attacks and prior knowledge of the training details of the model. Extensive experiments demonstrated that Model X-ray has outstanding effectiveness and efficiency across diverse backdoor attacks, datasets, and architectures. Besides, ablation studies on hyperparameters and more attack strategies and discussions are also provided.
Deep Learning, Backdoor Detection, Decision Boundary
$\dagger$ The corresponding author. E-mail: zhang_jie@cfar.a-star.edu.sg
journalyear: 2024copyright: rightsretainedconference: Proceedings of the 32nd ACM International Conference on Multimedia; October 28-November 1, 2024; Melbourne, VIC, Australiabooktitle: Proceedings of the 32nd ACM International Conference on Multimedia (MM ’24), October 28-November 1, 2024, Melbourne, VIC, Australiadoi: 10.1145/3664647.3681075isbn: 979-8-4007-0686-8/24/10ccs: Security and privacyccs: Computing methodologies Machine learning
1. Introduction
@@
Figure 1. Comparison of the decision boundaries between the clean model and the backdoored model (taking BadNets (Gu et al., 2017) as an example, and the target label is “airplane”) on the CIFAR-10 dataset.
@@
Despite the remarkable success of DNNs, recent studies (Gu et al., 2017; Li et al., 2021a; Zeng et al., 2021; Wang et al., 2022; Liu et al., 2018; Doan et al., 2021; Bagdasaryan and Shmatikov, 2021) have unveiled a significant security vulnerability for DNNs against backdoor attacks, which can contaminate DNNs, enabling them to operate normally on clean inputs but manipulate predictions when specific patterns (i.e., “trigger”) occur. Backdoor attacks primarily fall into two categories: data-poisoning attacks (such as BadNets (Gu et al., 2017), SSBA (Li et al., 2021a), Low Frequency (Zeng et al., 2021), and BPP (Wang et al., 2022)) and model-modification attacks (such as TrojanNN (Liu et al., 2018), LIRA (Doan et al., 2021), and Blind (Bagdasaryan and Shmatikov, 2021)). These attacks pose a substantial threat to safety-critical and security-sensitive applications of DNNs, including but not limited to face recognition (Parkhi et al., 2015), biomedical diagnosis (Esteva et al., 2017), and autonomous driving (Redmon et al., 2016). To mitigate the threat of backdoor attacks, numerous defense methods are emerging to establish a comprehensive pipeline for backdoor defense. This pipeline can be applied at various stages, including the training, post-training, and deployment stages (refer to Fig. 2).
@@
Backdoor defense during both the training and deployment stages (Chen et al., 2018; Tran et al., 2018; Gao et al., 2019; Zeng et al., 2021; Liu et al., 2023) typically necessitates access to training data or inference data. In this paper, we consider the more practical post-training scenario, where the defender aims to evaluate whether a trained model (e.g., Model Zoo that provides pre-trained models (zoo, 2018)) has been compromised by backdoor attacks, when and many post-training defenses assume the defender independently possesses a small set of clean, legitimate samples. However, current post-training detection methods hold too strong assumptions that the defender has knowledge of the attack information, the logit output from the model (Chen et al., 2019; Xu et al., 2021), and knowledge of the model parameters (Wang et al., 2019; Liu et al., 2019; Fu et al., 2023; Wang et al., 2023), limiting their implement in practical scenarios.
@@
Fortunately, recent work by (Somepalli et al., 2022) has demonstrated that we can visualize the model’s decision boundary solely using prediction labels. Leveraging this technique, we have identified a discernible distinction between the decision boundaries of the clean model and the backdoored model. As illustrated in Fig. 1, we use BadNets (Gu et al., 2017) as an example of backdoor attacks. We observe that the decision boundaries of backdoor models exhibit a noticeable reduction in the regions dominated by three clean samples, and significant surrounding area are dominated by the attack target label. Importantly, this phenomenon is applicable across various backdoor attacks on different datasets (see Fig. 4). That is to say, we can leverage the phenomenon of anomalous decision boundaries to distinguish backdoored models. As claimed in previous work (Wang et al., 2019), backdoor attacks build a shortcut leading to the target label, which we explain cause the above encircling phenomena. Besides, trigger samples are more robust against distortions (Rajabi et al., 2023), causing the large regions than that of clean samples. In a nutshell, the visualized 2D decision boundary can be served as an illustration for these conjectures.
@@
Based on the intriguing phenomenon, drawing an analogy to X-rays in disease diagnosis, we propose Model X-ray  as a novel backdoor detection approach through the analysis of illustrated 2D decision boundaries. Specifically, we designate two metrics to evaluate the degree of the closeness of the decision boundary: 1) Rényi Entropy (RE) (Rényi, 1961) calculated on the probability distribution of each prediction area and 2) Areas Dominated by triple samples (ATS), e.g., the total areas of “frog”, “ship”, and “dog” in the Fig. 1. Furthermore, if only one label is infected, we can determine the target label by the prediction of the largest area of the decision boundary, e.g., the target label is “airplane” in the right of Fig. 1. In other words, Model X-ray can not only identify backdoored models but also determine the target attacked label under all-to-one attacks. Importantly, Model X-ray accomplishes this only by the predicted hard labels of clean inputs from the model, regardless of any assumptions about attacks such as the trigger patterns and training details. The visualized 2D decision boundary offers a novel perspective to understand the behavior of the model, providing both visualization and interpretability. Through analysis of the decision boundary, Model X-ray can function as a lightweight diagnostic scanning tool, complementing other defense methods and aiding in defense pipelines. Extensive experiments demonstrate that Model X-ray performs better than current methods across various backdoor attacks, datasets, and model architectures. In addition, some ablation studies and discussions are also provided.
@@
  • We present a noteworthy observation: there exists a distinction between clean models and backdoored models as visualized through 2D decision boundaries (Somepalli et al., 2022). And we have visualized the decision boundaries of various backdoor attack models.
@@
  • We propose Model X-ray  which detects the backdoored model solely by predicted hard labels of clean inputs from the model, regardless of any assumptions about backdoor attacks. Besides, Model X-ray can determine the target attacked label if the attack is all-to-one attack.
  • We evaluate Model X-ray using four combinations of datasets and architectures (including CNNs and ViTs) and seven backdoor attacks with diverse trigger types. Additionally, we assess Model X-ray on different backdoor attack label mapping strategies and varying poisoning rates. Extensive experiments demonstrate that Model X-ray has outstanding effectiveness and efficiency.
2. Related Work
2.1. Backdoor Attacks
(1) $\begin{array}[]{r}\theta=\arg\min_{\theta}\mathbb{E}_{(x,y)\sim\mathcal{D}}\mathcal{L}(\hat{M}(x;\theta),y)\\ +\mathbb{E}_{\left(\hat{x},y_{t}\right)\sim\mathcal{\hat{D}}}\mathcal{L}(\hat{M}(\hat{x};\theta),y_{t}),\end{array}$
where $\mathcal{D}$ and $\mathcal{\hat{D}}$ denote the benign samples and trigger samples, respectively. $\mathcal{L}$ denotes the loss function, e.g., cross-entropy loss. The infected model functions normally on benign samples but yields a specific target prediction $y_{t}$ when presented with trigger samples $\hat{x}$. Backdoor attacks can be achieved by data poisoning and model modification, and we briefly introduce some related methods below.
Data poisoning-based backdoor attacks primarily revolve around crafting trigger samples. Notably, BadNets (Gu et al., 2017) was a pioneering work that highlighted vulnerabilities of DNNs by employing visible squares as triggers. Afterward, various other visible trigger techniques have been explored: Blended (Chen et al., 2017) employs image blending to create trigger patterns, SIG (Barni et al., 2019) utilizes sinusoidal strips as triggers, and Low Frequency (LF) (Zeng et al., 2021) explores triggers in the frequency domain. Simultaneously, other research endeavors focus on achieving imperceptibility of the trigger patterns, including BPP (Wang et al., 2022) based on image quantization and dithering, WaNet (Nguyen and Tran, 2021) founded on image warping, and SSBA (Li et al., 2021a) achieved by image steganography. During the training stage, the attacker can leverage different poisoning ratios to balance the attack ability and performance degradation.
Apart from data poisoning-based attacks, there are some backdoor attacks that employ model modification techniques. TrojanNN (Liu et al., 2018) first proposes to optimize the trigger to ensure that the crucial neurons can attain their maximum values, LIRA (Doan et al., 2021) formulates malicious function as a non-convex, constrained optimization problem to learn invisible triggers through a two-stage stochastic optimization procedure, and Blind (Bagdasaryan and Shmatikov, 2021) modifies the training loss function to enable the model to learn the malicious function.
2.2. Backdoor Defenses
Figure 2. The pipeline of the backdoor defense.
Comparably, post-training backdoor detection is model-level detection. Neural Cleanse (Wang et al., 2019) is the first post-training detection through anomaly analysis on the reversed trigger patterns. However, it requires access to the model’s inner information like parameters and gradients, which is also the limitation of other subsequent methods (Wang et al., 2019; Liu et al., 2019; Wang et al., 2020; Fu et al., 2023; Wang et al., 2023). Differently, detection work in black-box scenarios is extremely challenging (Chen et al., 2019; Xu et al., 2021; Guo et al., 2021; Dong et al., 2021), e.g. MNTD trains a meta-classifier based on features extracted from a large set of shadow models. However, its success heavily relies on the generalization capability of the attack settings from the shadow models to the actual backdoored models. Besides, it requires the soft label generated by the target model. MM-BD (Wang et al., 2023) leverags maximum margin statistics of each class and unsupervised anomaly detection on classifier output landscapes.
2.3. Decision Boundary of Deep Neural Networks
Most previous works depict decision boundaries by adversarial samples (He et al., 2018; Khoury and Hadfield-Menell, 2018) or sensitive samples (He et al., 2019). These methods are pivotal in identifying and understanding the contours of decision boundaries, as adversarial and sensitive samples are typically positioned along these critical junctures in the model’s decision-making process. However, obtaining these special samples requires access to the target model. Fortunately, Zhang et al. (Zhang et al., 2017) find that decision boundaries not only manifest near the data manifold but also within the convex hull created by pairs of data points.
Leveraging this understanding, Somepalli et al. (Somepalli et al., 2022) introduce an innovative approach that utilizes only clean samples to map out the decision boundary to investigate reproducibility and double descent. Their method, which results in a 2D map, offers an intuitive and accessible means of visualizing decision boundaries. In this paper, we utilize this technique to detect backdoored models.
3. Preliminaries
3.1. Recap of the Decision Boundary in (Somepalli et al., 2022)
After representing the space, we can calculate the bounds on the X-axis and the Y-axis, extended by a factor of $\eta$ in both the positive and negative directions along the corresponding axes, serving as a means to control the expansion range of the coordinate system. In the previous work (Somepalli et al., 2022), $\eta$ is set as 1 to investigate reproducibility, while we set $\eta$ as 5 to obtain a wider range of the decision boundary (see Fig. 3). Moreover, we can also determine density $S$ by constructing the set of points with a quantity of $S^{2}$ within the bounded range of the coordinate system using a grid generation method. Larger $S$ means higher resolution. With $S^{2}$ points, we can conduct the reverse process to get their tensor presentation, which can be fed to the model to fetch the corresponding prediction. We adopt different colors for different predictions to get the final 2D decision boundary.
Figure 3. Visual examples of the decision boundary used in (Somepalli et al., 2022) (left) and in this paper (right).
3.2. Threat Model
In practice, acccess to training or inference sets is unavailable due to data privacy, ownership, and availability constraints. Therefore, in this paper, we only consider the post-training scenario detection.
While many post-training detection defenses typically have access to either the model’s weights(Wang et al., 2019; Liu et al., 2019; Wang et al., 2020; Fu et al., 2023; Wang et al., 2023) or the model’s logit output(Xu et al., 2021) for evaluation, our approach goes a step further by restricting access to the model. We only assume that the defender has the capability to independently gather a small set of clean data samples that cover all classes within the domain, a prerequisite upon which most post-training detectors depend. Moreover, we only need the hard label predictions of the target model.
4. Method
In this section, we first provide an intriguing observation on the decision boundary of clean models and backdoor models. Based on this, we designate two strategies for backdoor detection via the decision boundary. Finally, we showcase that we can determine the target attacked label, if only one single label is infected.
Figure 4. Visual examples of decision boundaries of the clean model and different backdoored models on CIFAR-10 and ImageNet-10.
4.1. An Intriguing Observation
As shown in Fig. 4, we provide the decision boundary of the clean model and different backdoored models (infected by BadNets (Gu et al., 2017), SSBA (Li et al., 2021a), LF (Zeng et al., 2021), BPP (Wang et al., 2022), TrojanNN (Liu et al., 2018), LIRA (Doan et al., 2021), and Blind (Bagdasaryan and Shmatikov, 2021)) on CIFAR-10 and ImageNet-10 dataset. We observe two key phenomena in backdoored models: (1) the noticeable shrinking of areas dominated by clean samples and (2) the significant increase in surrounding areas dominated by target labels. More visualized decision boundaries can be found in the supplementary material.
We explain this phenomenon may be the shortcut effect caused by backdoor attacks. In essence, clean models can still preserve the robustness of predicted labels when applying a linear transformation to samples in a considerably large magnitude. On the contrary, the presence of shortcuts to the target label in backdoor models results in changes in the predicted label when applying a minor linear transformation to samples, typically leading to the target attacked label. The shortcuts leading to the target attacked label in the backdoor model has been confirmed in previous research, that is, through optimization methods, smaller perturbations can be found to cause other labels to be misclassified as target labels (Wang et al., 2019). Afterward, Rajabi et al. (Rajabi et al., 2023) quantifies this effect by introducing the concept of a certified radius (Cohen et al., 2019), which estimates the distance to a decision boundary by perturbing samples with Gaussian noise with a predetermined mean and variance. Notably, trigger samples are observed to be relatively farther from a decision boundary compared to clean samples, which can support why the large region is dominated by injected prediction.
Figure 5. Illustration on calculation of RE and ATS.
4.2. Two Strategies for Backdoor Detection via the Decision Boundary
Based on two key phenomena, we have developed two corresponding strategies that focus on the decision areas dominated by clean samples and the concentration of label distribution, namely, based on Rényi Entropy (RE) and Areas dominated by Triple Samples (ATS).
4.2.1. Backdoor Detection based on Rényi Entropy
With the technique mentioned above, we can plot $N$ decision boundaries $\mathbf{B}=\{\mathcal{B}_{1},...,\mathcal{B}_{k},...,\mathcal{B}_{N}\}$, where $\mathcal{B}_{k}$ is plotted along the plane spanned by triple samples $T_{k}=(x_{1},x_{2},x_{3})_{k}$. Specifically, let $S_{k}=\{x_{ij}|(i,j)\in\mathcal{B}_{k}\}$ be the set of points in the $\mathcal{B}_{i}$, where $(i,j)$ is the coordinations of $x$ in $\mathcal{B}_{k}$. Then, we feed $S_{k}=\{x_{ij}|(i,j)\in\mathcal{B}_{k}\}$ to the target model $M$ to obtain the corresponding hard labels $L_{k}=\{l_{ij}|(i,j)\in\mathcal{B}_{k}\}$, which are further used to obtain the final colorful decision boundary $\mathcal{B}_{k}$ for evaluation.
(2) $\small p_{m}=\frac{A(l_{m})}{A(\mathcal{B}_{k})},$
@@ 4.2.1. Backdoor Detection based on Rényi Entropy
(3) $\small RE(\mathcal{P}_{k})=H_{\alpha}(\mathcal{P}_{k})=\frac{1}{1-\alpha}\log\left(\sum_{m=1}^{n}p_{m}^{\alpha}\right),$
4.2.2. Backdoor Detection based on Areas dominated by Triple Samples
(4) $\small ATS(\mathcal{B}_{k})={\frac{A(T_{k})}{A(\mathcal{B}_{k})}}={\frac{\sum_{x\in{(x_{1},x_{2},x_{3})}}{A(x)}}{A(\mathcal{B}_{k})}},$
@@ 4.2.2. Backdoor Detection based on Areas dominated by Triple Samples
Figure 6. The label probability distribution within decision boundaries of clean and backdoor models on CIFAR-10 and ImageNet-10, both of whose infected labels are 0.
4.3. Determine the Target Label
5. Experiment
5.1. Experimental Settings
Datasets and Architectures. The datasets include CIFAR-10 (Krizhevsky et al., 2009), CIFAR-100 (Krizhevsky et al., 2009), GTSRB (Houben et al., 2013), and ImageNet-10 (imt, 2020), a subset of ten classes from ImageNet (Deng et al., 2009). Besides, we employ four different architectures: PreActResNet-18 (He et al., 2016), MobileNet-V3-Large (Howard et al., 2019), PreActResNet-34 (He et al., 2016), and ViT-B-16 (Dosovitskiy et al., 2020). These architectures encompass both Convolutional Neural Networks (CNNs) and Vision Transformers (ViTs) and span across various network sizes, including small, medium, and large networks.
Implementation Details. For the model to be evaluated, we plot decision boundaries by random samples triplet with expansion factor $\eta=5$ and density $S=100$, number of plots $N=20$. For the attack baselines, we evaluate our method against seven backdoor attacks, including BadNets (Gu et al., 2017), SSBA (Li et al., 2021a), LF (Zeng et al., 2021), BPP (Wang et al., 2022), TrojanNN (Liu et al., 2018), LIRA (Doan et al., 2021), and Blind (Bagdasaryan and Shmatikov, 2021). We follow an open-sourced backdoor benchmark BackdoorBench (Wu et al., 2022) for the training settings of these attacks and conduct all-to-one attacks by default. As shown in Table 1, the attacks in our experiments include both data poisoning-based attacks and model modification-based attacks, which contain diverse and complex trigger pattern types. In this paper, our focus is on post-training backdoor detection. We compare our approach with three post-training detection methods: Neural Cleanse (Wang et al., 2019), MNTD (Xu et al., 2021), and MM-BD (Wang et al., 2023). We utilize their official implementations (mnt, 2020; mmb, 2023) or implementations available in open-source benchmarks(Bac, 2023).
Table 1. The backdoor attacks involved in our evaluations have covered diverse trigger patterns.